Determine whether the equation represents y as a function x
A. x²+y²=4
B. x²+y²-2x-4y+1=0

Found 3 solutions by josgarithmetic, ikleyn, MathTherapy:
Answer by josgarithmetic(39800)

Neither.

x^2 and y^2 in both.

Answer by ikleyn(53765)

.
Determine whether the equation represents y as a function x
A. x²+y²=4
B. x²+y²-2x-4y+1=0
~~~~~~~~~~~~~~~~~~~~~~

Both cases have the same answer: NO.

Because in each case each value of "x" in the domain area determines TWO values of "y".

In order for "y" be the function of "x", each value of "x" must determine only ONE value of "y".


Answer by MathTherapy(10810)


Determine whether the equation represents y as a function x
A. x²+y²=4
B. x²+y²-2x-4y+1=0
Both equations represent CIRCLES, and circles are NOT functions as they fail the vertical line test.